slove the equations by substitution method
x+3y=10;2x+y=5

Answer by Maths68(1474)

x+3y=10..........(1)
2x+y=5...........(2)
Take (1)
x+3y=10
x=10-3y
Substitute the value of x in (2), we have
2(10-3y)+y=5
20-6y+y=5
20-5y=5
-5y=5-20
-5y=-15
Divide by (-5)both sides of above equation
-5y/-5=-15/-5
y=3
Plug in the value of y in either equations (1) or (2)
We take (1)
x+3y=10..........(1)
x+3(3)=10
x+9=10
x=10-9
x=1
Solution set (1, 3)
Check
Put the value of x and y in (2)
2x+y=5...........(2)
2(1)+3=5
2+3=5
5=5